A ball is realeased from a tower of height 40 m with velocity u. it reaches to the ground in 4sec. find u

Grade:9

1 Answers

kittu10
7 years ago
S=Ut+1/2gt2 here s=40, u=?, t=4. So, 40=4u+1/2×10×16. Solving this we get, u=(-10)msec-1. You may think about the neg sign. It states that the ball was thrown up wards first. It reached a height of 5 m upwards. You can calculate from (v)2=(u)2-2gs. Where v=0. You will see that s=5. Time required=1sec.So now total height is 40+5=45m. At this time v=0. So calculate the time by putting s=45, u=0, in s=ut+1/2g (t)2. Find out t. T=3. That means total time=1+3=4sec. Hope this helps. Inform me if I'm wrong. That will help me too!